OpenStudy (waleed_imtiaz):

x% of x is same as 20% of : A) x^2/20 B) x^2/10 c) x^2/100 D) 20/x^2 please show me the opinion and logic.......

OpenStudy (anonymous):

x% of x is basically x^2/100. Now, let x^2/100 be 20% of y, where y is the answer we're looking for. We have: x^2/100 = 20% * y = y/5, which gives us y = x^2/20. Answer: A

OpenStudy (anonymous):

\[\large \frac{x}{100} \times x =\frac{20}{100} \times y \implies x^2 = 20y \implies y = \frac{x^2}{20}\]

OpenStudy (waleed_imtiaz):

oh.... yes.. I got it.... thanku guyz